Operations Report – January 26th

MDRS Crew 220

Operations Report

Crew 220 Operations Report 26JAN2020
Sol: 7

Name of person filing report: Matt Wise

Non-nominal systems: Primary Generator, Water Heater

Generator: Primary is inoperable. Rental was run-the following information is for the rental.

Hours run: 11hr 40min
From what time last night: 17:20

To what time this morning: 07:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 91%

Diesel Reading – Below 1/2 (approx 48%)

Station Propane Reading- 85%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0148290,8

Water (static tank) (gal) – 241

Static to loft Pump used – Yes

Water in Green Hab (gal): 163.4

Water in Science Dome (gal): 0

Toilet Tank Emptied – Yes

Deimos rover used: Still in the workshop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: No

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – Yes

Starting Hours: 127.9

Beginning Charge: 100%

Ending Hours: 128

Ending Charge: 89%

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Starting Hours: 133.9

Beginning Charge: 100%

Ending Hours: 134

Ending Charge: 85%

Notes on rovers: Opportunity and Deimos off-site for maintenance. There were no EVAs today, so the rovers were not used.

ATVs Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today: 0

Notes on ATVs: None.

Habcar used and why, where? No

CrewCar used and why, where? No

General Notes and Comments: None.

Summary of the internet: Nominal

Summary of Suit and Radios: Nominal

Summary of Hab: Main lights in hab lower deck dim and flickering

Summary of Science Dome operations: Nominal

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and Safety issues: Please refer to HSO report

Questions, concerns, and requests to Mission Support: None. Thanks!

Operations Report – January 26th

MDRS Crew 220

Operations Report

Crew 220 Operations Report 25JAN2020

Sol: 7

Name of person filing report: Matt Wise

Non-nominal systems: Primary Generator, Water Heater

Generator: Primary is inoperable. Rental was run-the following information is for the rental.

Hours run: 11hr 40min
From what time last night: 17:20

To what time this morning: 07:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 91%

Diesel Reading – Below 1/2 (approx 48%)

Station Propane Reading- 85%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0148290,8

Water (static tank) (gal) – 241

Static to loft Pump used – Yes

Water in Green Hab (gal): 163.4

Water in Science Dome (gal): 0

Toilet Tank Emptied – Yes

Deimos rover used: Still in the workshop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: No

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – Yes

Starting Hours: 127.9

Beginning Charge: 100%

Ending Hours: 128

Ending Charge: 89%

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

Starting Hours: 133.9

Beginning Charge: 100%

Ending Hours: 134

Ending Charge: 85%

Notes on rovers: Opportunity and Deimos off-site for maintenance. There were no EVAs today, so the rovers were not used.

ATVs Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today: 0

Notes on ATVs: None.

Habcar used and why, where? No

CrewCar used and why, where? No

General Notes and Comments: None.

Summary of the internet: Nominal

Summary of Suit and Radios: Nominal

Summary of Hab: Main lights in hab lower deck dim and flickering

Summary of Science Dome operations: Nominal

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and Safety issues: Please refer to HSO report

Questions, concerns, and requests to Mission Support: None. Thanks!

Sol Summary – January 26th

Crew 220 (MDRS) Sol Summary 22Jan2020

Sol: 7

Summary Title: Soulful Sol

Author’s name: Connie Delisle

Mission Status: Marsvellously Nominal

Sol Activity Summary:

The day started with a joint briefing and breakfast between the two Crews after MAU Crew came over to the MDRS Hab. The nature of this Mission intentionally provides opportunity for joint EVAs and learning opportunities in order to observe and discuss the nature of small-larger group human factor interaction. Today’s morning activities included medical training on the METHANE procedure, additional drilling on the P-MARCH-P, and a training session on CRP and bagging (resuscitation). Crew members reported learning a great deal and appreciating the hands-on practice and drilling. Today’s activities with a smaller Crew followed the planned schedule and timing more closely than in previous days. In addition, the Crew reported being more relaxed and feeling less hurried between planned events.

Look Ahead Plan: Sol 8 includes a full schedule, starting at 07:00 for breakfast and MDRS-MAU Commander briefing on the day’s events. Following will be a Medical EVA near the MDRS Hab, serving as the initial field test of the “Golden Bubble.” This ground-breaking technology enabled innovation represents the culmination of several months of intense design work. The evening events will include: social science research (nutrition study, psycho-social study on willingness to go to Mars and a “Grand Rounds” technical presentation Sensory Illusion Cognitive Bias and False Memory from one of the Crew members who is a medical doctor.

Weather: Tonight, cloudy skies with a lowest temperature around -2°C, with a 40% chance of snow. The average highest temperature during the day until Saturday is 5°C, and the lowest during the night would be around -4°C.

Crew Physical Status: The Crew has been working hard and is somewhat fatigued. However, they were bolstered by the wonderful day on Mars.

EVA: One afternoon EVA was conducted, starting at MDRS to MAU Station, with a return to MDRS at 17:30.

EVA Report – January 26th

Crew 220 EVA Report 26Jan2020

EVA # 1

Author: Johannes Svensoy, MD

Purpose of EVA: Moving personnel and equipment between MDRS and MAU Stations.

Start time: 1500

End time: 1700

Narrative:

PRE-EVA: Midday was used to pack and to go through equipment. All EVA Crew were hydrated and fed before leaving for EVA. The terrain here on Mars is muddy, however with no addition precipitation it is drying up some places.

EVA:

The airlock was loaded with personnel and equipment before depressurization, which was done without problems. Spirit and Curiosity rovers were used to move equipment over to MAU Station. After returning the rovers, the crewmembers walked over to MAU station doing some very nice terrain observations and photos on the way.

Destination: MAU Station

Coordinates: 4250050,518500

Participants: HabComm: Connie. EVA Crew: Johannes, Shawna, and Morgan.

Road(s) and routes per MDRS Map: Entrance road out to Cow Dung Road. Cow Dung Road South to MAU station.

Mode of travel: EVA with rover to move equipment.

Support Requested: No support requested during EVA.

Operations Report – January 26th

Crew 220 Operations Report 26JAN2020

Sol: 7

Name of person filing report: Johannes Svensoy, MD

Non-nominal systems: N/A

Generator: Not Run

Hours run: 0

From what time last night: N/A

To what time this morning: N/A

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): n/a

Diesel Reading – N/A

Station Propane Reading- 2%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – N/A

Water Meter (units)- N/A

Water (static tank) (gal) – N/A

Static to loft Pump used – N/A

Water in Green Hab (gal): N/A

Water in Science Dome (gal): N/A

Toilet Tank Emptied – no

Deimos rover used: N/A

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – Between MDRS and MAU Station.

Starting Hours: 127,9

Beginning Charge: 100%

Ending Hours: 128.0

Ending Charge: 89%

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Between MDRS and MAU Station.

Starting Hours: 133,9

Beginning Charge: 100%

Ending Hours: 134

Ending Charge: 85%

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where?

CrewCar used and why, where?

General Notes and Comments:

Summary of the internet: N/A

Summary of Suit and Radios: Nominal

Summary of Hab: Structural integrity nominal. Floors becoming dirty due to mud. Will be re-enforcing with cardboard and tarps

Summary of Science Dome operations: Temperatures are hovering around 7°C during the day.

Summary of RAM operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: None

Eva Report – January 26th

Crew 220 EVA Report 26Jan2020 (MAU Station)

EVA # 1

Author: Johannes Svensoy, MD

Purpose of EVA: Moving personnel and equipment between MDRS and MAU Stations.

Start time: 1500

End time: 1700

Narrative: Incoming MDRS crew members helped the outgoing MDRS crew move their equipment over to MAU Station. The Spirit and Curiosity rovers were used to move equipment over to MAU Station. After returning the rovers, the crewmembers who are now going to stay at MAU Station walked over to MAU station to settle in. It was a nice walk between the stations taking in the beautiful Martian terrain along the way.

Destination: MDRS Station and MAU Station.

Coordinates (use UTM NAD27 CONUS): 4250050,518500

Participants: HabComm: Connie (MDRS). EVA Crew: Johannes, Shawna, Morgan.

Road(s) and routes per MDRS Map: Entrance road out to Cow Dung Road. Cow Dung Road South to MAU station.

Mode of travel: EVA with Rovers to move equipment.

Support Requested: No support requested during EVA.

Commander Report – January 26th

Summary Title: Sunday’s Trinity

Author’s name: Connie Delisle, Commander
Mission Status: Nominal
Sol Activity Summary: Along with transfer for personnel this morning came the transfer of Command of MDRS Station from Commander Shawna to Commander Connie; and at MAU, from Commander John to Commander Shawna. This first-ever MAU-MDRS Station-to-Station Mission purposefully includes command transfers to test differences between governance (leadership models). Several of the MDRS Crew will have coordinated three “departures” by the end of this day: Several Crew were quarantined; a joint MAU-MDRS EVA was done to the MAU Station using the Rovers (Spirit and Curiosity); and a return of the MAU Crew to its Station. We are finding synergy in “threes”. The morning activities at MRDS Station included medical training for non-medical professionals. The informative, hands on training session was co-presented by Dr. Johannes Svenszoy and MD Dr. Shawna Pandya, MD. The first training block covered the procedure for applying the METHANE procedure:[Major Incident, Exact Location, Type of incident, Hazards involved, Access to Site Number of Causalities, and Emergency Services Needed]. This session was followed by Dr’s Pandya and Svensoy reviewing the procedure for conducting P-MARCH-P [Patient, Personnel, Permission, Personal Protective Gear; Massive Hemorrhage and C-Spine; Airway; Respirations; Circulation and Head/Hypothermia. After lunch the Doctors resumed the training session, reviewing CPR and bagging (resuscitation) procedures inclusive of hands on practice using the technique.

Team lead Commander Shawna managed gracefully through some communications technology issues at the start of the afternoon EVA to the MAU station. Keeping safety first, the Crew successfully completed the EVA, capturing photos and returned to MDRS at 17:45. Dr. Pandya then completed the EVA wellness study which entails collecting biometrics data.

Look Ahead Plan: Tomorrow’s morning and afternoon Medical EVAs will consist of testing an innovation entitled the “Golden Bubble”. The Bubble is a prototype of a first-of-its kind pressured Medivac transport System.
Anomalies in work: NA
Weather: -3C low / 3C max, 0C average, 40% chance of precipitation tomorrow
Crew Physical Status: Excellent mental focus, with positive feedback about the high quality medical training sessions.
EVA: 1 AM MDRS – MAU joint EVA performed today as above

Reports to be filed:
Commander Report
Sol Summary
Photo Report
Operations Report
GreenHab Report
EVA Report
EVA Request
Support Requested: None Required

GreenHab Report – January 26th

Crew 220 GreenHab Report 26-JAN-2020 GreenHab Officers: Marufa Bhuiyan and Morgan Kainu

Environmental control: Heating GreenHab door shade cloth (40%) (30%) on.

Average temperatures: 19.25 °C 08:56 Floor Unit: 17 °C

10:44 Floor Unit: 19 °C

12:16 Floor Unit: 22°C

17:55 Floor Unit: 19°C

Hours of supplemental light: 19:30 to 00:00

Daily water usage for crops: 13.2 gallons

Daily water usage for research and/or other purposes: NA

Water in Blue Tank 163.04 gallons

Times(s) of watering for crops: 3 08:56 10:45 18:09

Changes to crops: Pumpkin flowers are blooming. Planted 1g sweet pepper (California Wonder), 1g spinach lakeside.

Narrative: Plants seem to be maintaining well and enjoying their feedings and sunlight.

Harvest: Crops picked for cooking: 1g basil, 3g cilantro, 2g acorn pumpkin leaves

Support/supplies needed: MDRS Seed Bank lid doesn’t close properly; it may affect the seeds.

Sol Summary – January 26th

Shawna
Commander,
MAU Crew 220

MDRS Crew 220 Sol 7 Summary 26Jan2020

Summary Title: Olympus MAUns Approacheth
Author’s name: Shawna Pandya, Commander
Mission Status: MAUminal
Sol Activity Summary: The day began with a skeleton crew, as several members are in quarantine as part of our simulation, so we are down to 6 between MAU-MDRS station, and look forward to working, growing and bonding together as a small group over the next few sols. We spent the morning going over additional medical drills and training in anticipation of our medical EVAs over the next few sols. Later on, Commander Pandya and XO Svensoy moved out to MAU station to take command of operations there, while acting Commander Delisle stepped up to run the MDRS operations. The days ahead will be full, and we all look forward to them. Our sol tonight will end with a "Grand Rounds" academic activity wherein one of our officers will teach the crew about an area of their passion, along with a crew round-table discussion of emergency and medical kit planning – and a little fun and bonding, of course.
Look ahead plan: Tomorrow, the crew will put their medical learning and drilling into practice over the course of a morning recap session, followed by two medical EVAs in front of the hab to simulate these skills in an EVA setting.
Weather: -3C low / 2C max, 0C average, 40% chance of precipitation tomorrow
Crew Physical Status: Starry-eyed, nominal
EVA: 1 PM EVA as old MDRS crew moved to MAU station
Reports to be filed:
Sol Summary
Commander’s Report (sent)
Ops/Engineering Report
GreenHab Report
Photo Report
EVA Request: To be submitted.
Support Requested: Just moral support, please.